OpenStudy (anonymous):

Choose the correct simplification of the expression 2 b over c all to the power of 3. 2 times b to the 3rd power all over c to the third power 8 times b to the third power all over c to the third power 2 times b to the third power all over c 2 times b all over c to the third power

OpenStudy (anonymous):

\[(\frac{2b}{c} )^3 = \frac{ (2b)^3 }{ c^3} = \frac{ 2^3 b^3 }{ c^3}\]

OpenStudy (anonymous):

\[2^3 = 2 \times 2 \times 2 = ?\]

OpenStudy (anonymous):

simplify for the final answer

OpenStudy (anonymous):

(2b / c)^3 (2b)^3 / c^3 2^3 b^3 / c^3

OpenStudy (anonymous):

its b

OpenStudy (anonymous):

Yes 2^3=8 so b is the answer :)

OpenStudy (anonymous):

can you help me with more?

OpenStudy (anonymous):

Of course

OpenStudy (anonymous):

i'll help with what i can!

OpenStudy (anonymous):

Choose the correct simplification of the expression a to the 6th power times b to the 3rd power all over a to the 4th power times b to the second power. a to the second power over b a10b5 a to the 10th power over b to the 5th power a2b

OpenStudy (anonymous):

\[\frac{ a ^{6} b ^{3}}{ a ^{4}b ^{2} }=a ^{2}b\]

OpenStudy (anonymous):

thank you I have one more!

OpenStudy (anonymous):

Part 1: Show the expanded form and simplification of x to the 6th power over x to the second power. Explain in your own words how you can simplify x to the 6th power over x to the second power without having to write the expanded form. Part 2: Create your own fraction with like bases, coefficients, and show its simplification.

OpenStudy (anonymous):

\[\frac{ (x)(x)(x)(x) }{ (x)(x) }=(x)(x)=x ^{2}\]

OpenStudy (anonymous):

you can simplify without writing the expanded form by subtracting the exponents

OpenStudy (anonymous):

\[\frac{ 16x ^{3}y ^{6} }{ 8xy ^{2} }=2x ^{3}y ^{4}\]

OpenStudy (anonymous):

thank you so much !